what is the graph using the slop-intercept method
2x+4y=10

Answer by RAY100(1637)

2x +4y = 10
.
4y = -2x +10
.
y= (-2/4)x +(10/4)
.
y= (-1/2)x +(5/2),,,,,since form is y=mx+b
.
slope (m) = -1/2,,,,,,,,and y intercept = 5/2
.
Plotting on x-y coordinates,,,(0,2.5),,,,(5,0),,,,connect points